OpenStudy (anonymous):

The sequence is geometric. What is the next term? 3125, 625, 125, 25, ... A. 20 B. 15 C. 5

OpenStudy (across):

Hint: This is \(5^x\), but backward.

sam (.sam.):

r=1/5 a=3125 Tn=ar^(n-1) T5=3125(1/5)^(5-1)=5